### v3

# Why Johnny Signs with Sigstore: Examining Tooling as a Factor in Software Signing Adoption in the Sigstore Ecosystem

Abstract
Software signing is the strongest method for ensuring the integrity and authenticity of components in a software supply chain. Although software signing has been recommended by regulation and industry consortia, practical adoption of software signing is generally limited. Previous studies have recognized tooling as a key factor in influencing software signing adoption and quality. However, most research has primarily focused on its user interface aspects, with little attention given to other usability considerations, such as case-specific user challenges and factors influencing tool adoption and changes.
To understand how software tools influence the practice and adoption of software signing, we study the formative usability of Sigstore, a modern and widely adopted software signing tool. We interviewed thirteen (13) experienced security practitioners to study the factors that influence the selection of a tool, the problems associated with the use of such tools, how practitioners’ software signing tools have evolved, and what drives this migration. To summarize our findings: (1) We highlight the various factors practitioner’s consider before adopting a software signing tool; (2) We highlight the problems and advantages associated with the current tooling choices of practitioners; and (3) We describe the evolution of tooling adoption of our sample population. Our findings provide the software signing tool development community with valuable insights to improve their design of software signing tools.
1 Introduction
The reuse of software components in modern software products results in the formation of complex software supply chains, leading to an increased attack platform for these products [39, 79, 67]. A key security strategy to mitigate this expanding attack surface is ensuring the authenticity of actors and integrity of artifacts by establishing provenance [54, 68]. Software signing is the strongest guarantee for establishing provenance [69] and has been extensively advocated by academia, industry consortia, and government regulations [81, 16, 14]. To effectively implement this process, software signing tools–such as OpenPGP [55], Sigstore [53], and Notary [57]–have been proposed and are currently used in practice to sign software artifacts.
Although tooling has significantly advanced various technological fields [71, 41, 3, 47], its impact on software signing must be carefully examined to maximize its benefits. While prior "Why Johnny Can’t Encrypt" inspired studies have explored the usability of signing tools, their focus has been limited primarily to user interface design [89, 73, 63]. Furthermore, their evaluation techniques rely primarily on summative usability assessments (ease of use), which have been criticized for being difficult to measure and for providing limited insights to inform tool enhancements [13, 48]. The applicability of these studies is also constrained by their narrow focus on message (and email) signing and their reliance on novice volunteer participants, which limits the real-world relevance of their findings for software signing.
Our work provides the first examination of the usability of tooling as a key factor influencing the adoption and practice of software signing. We conducted interviews with 13 experienced security practitioners to explore current software signing tools, with a focus on Sigstore as a case study. We apply a formative usability framework  [17], which emphasizes the iterative refinement of the studied tooling option, to develop our study instruments. Our study explores usability in two categories: users’ experience with the specified tooling option and the usability of the tooling options with respect to other software signing tools. Our study thus explores three key themes: the evolution and motivations behind software signing tool adoption, limitations and challenges encountered with currently adopted tooling, and the specific implementations of adopted tools.
Our results highlight the factors practitioners consider when adopting specific software signing tools, using a case study of a key software signing tool. This serves as formative feedback for tool designers, offering insights into the drivers of tool usage. We present these findings through a clear articulation of practitioners’ challenges and the merits of using particular tools. Additionally, we found that practitioners’ decisions to adopt a software signing tool are influenced by less commonly discussed factors, such as the tool’s community, relevant regulations and standards, and certain functionalities of modern Software signing tools.
In summary, we make the following contributions:
  1. 1.
    We report on the difficulties and advantages of using Sigstore, a next-generation software signing tool (section 5.1.1, section 5.1.2).
  2. 2.
    We discuss how practitioners’ software signing tools evolve over time and the factors that influence practitioners’ decisions to switch tooling options (section 5.2).
2 Background & Related Works
In this section, we discuss the foundations of software signing (section 2.1), usability studies in software signing (section 2.2), and the Sigstore ecosystem where we situate our work (section 2.3).
2.1 Software Signing
2.1.1 Signing Overview and Workflow
Software signing is a process that ensures the authorship and integrity of a software artifact by linking a cryptographic signature created by the software maintainer with their software artifact. Fundamentally, software signing is based on the public-key cryptography model  [69], and a traditional signing workflow follows a three-stage process of creation, distribution, and verification. We describe a typical signing workflow in Appendix C.
While the fundamental software signing process remains consistent across implementations, different signing tools and policies are shaped by the target platform, software artifacts, and root of trust. The choice of signing tools and the policies under which developers operate often depend on the operating system and distribution vendor. For example, Microsoft Authenticode [30] and Apple Xcode [33] provide platform-specific signing mechanisms, while package registries like PyPi [45] and Maven [1] enforce signing policies through dedicated tooling plugins. Developers also have a variety of software signing tools available to them, such as GNU Privacy Guard (GPG) [58], Signify [85], Sigstore [53], and Notary [57]. Each of these tools present different advantages and challenges, shaping how developers integrate signing into their workflows.
2.1.2 History and State-of-Art Tooling
Traditional Signing Tools: Many traditional signing tools, such as Git Commit Signing and GNU Privacy Guard are based on PGP. PGP (Pretty Good Privacy) is a proprietary encryption software originally developed by Phil Zimmerman in 1991. It is a public key cryptography-based protocol whose capabilities include encrypting files, creating and managing public and private keys, and creating digital signatures. Its open-source standard is OpenPGP [23, 55]. However, two main struggles associated with PGP-based software signing involve key management and establishing trust in public keys.
PGP relies on a "web of trust" model, where a signer’s public key is considered trustworthy based on endorsements from other trusted signers [23]. Each created public key is stored in a "key certificate" that specifies the public key, the name and/or email address associated with the public key, its creation date, and others who have signed the key. The signer is also responsible for secure management of their private key, preventing exposure, as the combination of their private key and passphrase enables the signing of artifacts [45].
State-of-art Signing Tools: Although software signing is widely recognized as an effective technique for mitigating software supply chain attacks, several obstacles hinder its adoption. Research highlights the prevalence of unsigned artifacts in software registries, emphasizing the need for improved signing practices [53, 90, 46]. This has led to the emergence of the next generation of software signing. Newman et al. introduced a new software signing platform in 2021 called Sigstore that addresses key adoption barriers, particularly key management, by leveraging keyless/ephemeral software signing techniques [53].
Keyless signing mitigates security risks associated with private key compromise. Instead of relying on a long-term private key, the signer uses a short-lived key that exists only for the duration of the signing process. This approach enhances security by preventing private key leaks while improving accessibility and usability [2]. As a result, many software registries are transitioning to this model to streamline verification and enforce stronger software integrity measures [15, 34]. While Sigstore’s development aims to solve the usability issues associated with traditional signing methodology, firsthand perspectives from practitioners using Sigstore will provide valuable insight into its actual usability and convenience.
2.2 Usability Studies in Software Signing
2.2.1 Usable Tools: Definition and Theory
Usable Tools: In basic terms, a software engineering tool assists developers in their software production process [27]. For software signing specifically, these tools are used to provide a method for which users can verify the integrity and authenticity of the code to be published, and developers use these tools to verify signatures of the software artifacts they may incorporate into their own projects. Deciding which software engineering tools to use is a critical question developers face when many options are available. A specific tool can affect a developer’s workflow efficiency, productivity [44], and the resulting quality of the work that is produced. This allows members of the computing industry to remain competitive by selecting the most suitable tools and environments that are conducive for the efficient production of software [27]. Additionally, tooling is important because it enables engineers to use software development practices they have not in the past [11].
Usability Evaluation: Studies assessing implementation strategies of products are typically framed as usability evaluations [49]. Usability evaluations are generally categorized into two forms: summative (validation or ease of use) and formative (diagnostic usability) [48]. Summative usability refers to the extent to which a system or product can be used to achieve a specified goal effectively, efficiently, and satisfactorily within a given context [35, 65]. In contrast, formative usability focuses on product assessments aimed at identifying areas for improvement [82]. Although summative usability evaluations are more commonly used, they are not designed to provide actionable feedback for enhancing the product [48]. Summative evaluations primarily measure metrics such as effectiveness, efficiency, and user satisfaction [35]. However, such metrics have been widely criticized for being subjective and difficult to quantify [48, 13].
In contrast, formative usability evaluations aim to diagnose user challenges and provide iterative feedback to enhance product functionality [82, 48]. These evaluations emphasize real-world usability issues and are more suited for refining a product’s design. While summative evaluations are more common, their limitations highlight the need for formative assessments when improving software signing tools.
2.2.2 Empirical Studies
Signing Tooling Usability Studies: The usability of digital signing tools has been widely studied, though often within the context of encryption rather than broader software signing applications. Whitten and Tygar’s seminal study, Why Johnny Can’t Encrypt [89], found that PGP 5.0’s user interface was too complex for non-expert users, making encryption and signing inaccessible. Sheng et al. [73] later observed that PGP 9’s interface had not significantly improved, particularly in key verification, encryption transparency, and digital signing usability. Other studies [10, 64, 60] have highlighted persistent usability issues, including the complexity of the public-key model.
Further research has explored the differences between manual and automated signing tools. Fahl et al. [19] and Ruoti et al. [62] found no significant usability differences between these approaches. Meanwhile, Atwater et al. [5] demonstrated that users preferred automated solutions based on the PGP standard. While these studies provide insights into encryption tools, their relevance to broader software signing workflows is limited.
These studies exhibit weaknesses in their scope, methodology, and the relevance of their findings to software signing. First, they primarily focus on evaluating the usability of signing tooling in the context of email or message encryption, limiting their application to software signing. Additionally, they limit their scope to user interface design, overlooking critical factors such as tooling adaptability and macroenvironment (e.g., tooling user communities). Furthermore, much of this research is outdated, reflecting decades-old technologies and failing to incorporate contemporary practitioner experiences. Their methodology predominantly employs summative usability assessments, which focus on ease of use [13]. These studies typically rely on novice volunteers to evaluate usability. In addition to the flaws of the summative evaluation approaches, the reliance on novice participants limits the applicability of their findings. Evaluating usability through inexperienced users does not yield specialized feedback that can effectively guide tool enhancements [42, 50].
2.2.3 Related Software Signing Research
The use of software signing as a security method to mitigate Software Supply Chain security has been studied in the literature. These studies has focussed on the adoption and practice of Software signing. Schorlemmer et al. [69] examine factors influencing software signing adoption through a framework of signing events. They identify tooling as a major factor affecting the quality of software signatures but focus primarily on the role of dedicated tooling within software registries, rather than its broader impact on engineering workflows. In our parent study, Kalu et al. [37], we provided valuable insights into the implementation and adoption frameworks of software signing, with a focus on broad software signing challenges faced in commercial settings. These challenges included issues of usability, compatibility, and integration. However, while we highlighted these broad challenges, our findings lacked the specificity needed to identify which tools are most impacted by these issues and under what conditions of use, limiting practitioners’ contextual understanding of these problems.
These prior studies acknowledge that improved tooling can drive the adoption of software signing. However, they fail to clearly define what constitutes an "improved tool" from the perspective of practitioners. While they emphasize broad usability factors, they overlook the specific technical and workflow-related requirements that practitioners consider when selecting tools for real-world use.
2.3 Sigstore
Due to the emergence of next-generation software signing and the growing adoption of Sigstore as a signing platform, we focus on the Sigstore tool specifically in our research.
Importance: Sigstore, launched in 2021, was developed to address critical challenges in software signing, including complex key management, usability barriers, and software supply chain security risks. Traditional signing tools, such as OpenPGP and enterprise certificate authorities, rely on long-term cryptographic key storage, which introduces risks of key compromise, revocation difficulties, and usability friction. In contrast, Sigstore simplifies the signing process by providing keyless, ephemeral authentication mechanisms. Within 13 months of its public release, Sigstore recorded 46 million artifact signatures  [29]. It has since gained significant adoption and support from open-source ecosystems  [84, 25, 31, 6, 86] and various organizations  [38, 76, 75, 77]. Given its growing adoption and integration in major software ecosystems, Sigstore represents a state-of-the-art software signing approach, making it a compelling case study for examining usability and security trade-offs.
Design Aspects Affecting Usability: One key challenge that is associated with traditional signing methods is key management, which is addressed by Sigstore. In conventional certificate authority-based signing, such as Windows Authenticode, a software artifact’s validity is linked to the certificate’s status—whether it is still valid, expired, or revoked [43, 70]. A revoked certificate can indicate compromise, but it may not be immediately clear when the compromise occurred [40]. This uncertainty allows some revoked certificates to be used for signing malware while still appearing valid to verification systems. Sigstore mitigates this risk by issuing short-lived certificates. These certificates significantly reduce the chances of key and certificate compromise since they automatically expire after a short duration, limiting their potential misuse.
Complexity: Sigstore uses a modular design of several components to implement software signing. As described by [53], Sigstore consists of eight key components: signers, verifiers, the artifact log, an OpenID Connect (OIDC) provider, a certificate authority, monitors, transparency logs, and package repositories. While Sigstore is designed to function as a cohesive system, these components can also be implemented individually based on specific use cases. The signing process starts with the signer using OpenID Connect (OIDC) to obtain a signing certificate. Unlike traditional methods requiring long-term key management, Sigstore associates a signer’s identity with short-lived tokens issued by services such as Google, GitHub, and Microsoft. Next, a software signing certificate is issued to the signer based on their verified identity. In this step, Fulcio serves as the certificate authority (CA), ensuring that the OIDC token was correctly issued to the requesting signer before granting the certificate. To monitor signing behavior, a verifier makes use of Fulcio’s log and the transparency log (Rekor) to verify for identity of the creator of the signatures, the certificate validity, and signature validity.
3 Knowledge Gaps & Research Questions
The existing literature on software signing, as outlined in  section 2, has examined its practice and adoption across different contexts, including private and open-source ecosystems. These studies highlight the critical role that the choice of a software signing tool plays in ensuring the quality of the software signing process.
Although certain aspects of software signing tools have been explored, such as the usability of the user interface design and functional exclusivity (dedication of tooling resources), the broader impact and role of software signing tools remain largely understudied. The findings of these works underscore that tooling is a multi-faceted and multi-layered factor. A general usability evaluation of these tooling options would significantly enhance their adaptability and facilitate improvements. Also, the existing research on software signing tooling primarily relies on experiments with novice users and quasi-experiments to investigate isolated aspects of tooling. Practitioner perspectives on the nuanced role of tooling in shaping the use and practice of software signing remain limited, particularly regarding how the usability of these tooling contributes to their adoption.
Usability Framework:
Figure 1: Usability Framework from Cresswell et al. [17]. Their framework assesses usability based on four factors: Technological (T), Social/Human (People), Organizational (O), and Macroenvironmental (M) factors. Our framework applies their framework to two concerns of software signing tool usage: (1) the tool itself and (2) previous signing experiences.
To examine the usability of software signing tooling, an examination of experienced practitioner experiences is necessary. As noted in section 2.2, the summative usability approach mostly is mostly hindered by its use of subjective measurement metrics. To effectively, answer usability in terms of how software signing tooling are used in practice, a formative usability framework is warranted. A formative framework addresses the experiences of regular users of a tool, rather than controlled experience of experimental subjects  [48].
Our formative usability evaluation framework, Figure 1, is based on Cresswell et al.’s [17] formative framework. Their framework identifies four key factors that influence the overall usability of a product or technology: Technological (T), Social/Human (People), Organizational (O), and Macroenvironmental (M) factors. Usability is determined by the interplay of these factors.
Our formative evaluation applies of their four-factor framework to two key areas: users’ experiences with the specified tooling option and the usability of the tooling option compared to previous software signing tools. We designed our study instrument to address these broad usability themes, emphasizing the identified usability factors.
Research Questions: Our research addresses the following questions in light of the positions outlined above within the context of Sigstore:
  • RQ1:
    What drawbacks and benefits of software signing tools influence their adoption in practice?
  • RQ2:
    What factors influence practitioners’ decisions to switch software signing tools over time?
4 Methodology
Figure 2: Study Methodology. We developed our interview protocol by reviewing Sigstore grey literature and the academic literature using our adopted formative usability framework. Finally, we analyzed our results using a thematic analysis informed by our usability evaluation framework (Figure 1).
We undertook a qualitative case study of the Sigstore software signing ecosystem through interviews of engineers. We applied a formative usability framework to identify the key factors that impact the adoption of Sigstore in practice.
Our methodology is illustrated in Figure 2. This study was approved by our Institutional Review Board (IRB).
4.1 Research Design and Rationale
4.1.1 Rationale
The case study is a research method used to investigate a phenomenon within its real-life context  [61, 91, 18, 24]. We selected the case study method for our research because it allows for an in-depth exploratory examination of our chosen case, enabling a deeper understanding of the usability factors that influence the adoption of a software signing tool in practice, rather than a broad but superficial cross-analysis  [59]. In selecting Sigstore, we adhered to the criteria outlined by Yin [91] and Crowe et al. [18]. We considered the intrinsic uniqueness of the unit (high), access to data (adequate), and risks of participation (low).
Our research objectives explore the impact of Software Tooling as a factor in the practice of software signing from the viewpoint of practitioners, with a particular focus on how the usability of software signing tools influences their adoption. We opted for semi-structured interviews [66] to effectively examine various aspects of tooling, including usability challenges, perceived benefits, and comparative experiences with other signing tools. This approach allows us to gather descriptive, open-ended, opinionated, and detailed responses, providing deeper insights into how usability considerations shape adoption decisions.
4.1.2 Instrument Design & Development
To develop our interview instrument, we relied on the usability framework (Figure 1) and various signing literature. We outline how these resources informed our instrument design:
  1. 1.
    Grey Literature: To derive our questions, we primarily examined Sigstore blogs [74], and Sigstore usage reports [8, 86, 31, 38, 76, 75, 77, 15]. These sources included reports on how various entities implemented Sigstore in their engineering processes, news about major updates, and other related Sigstore developments.
  2. 2.
    Background & Literature Review (section 2): We employed a snowball approach to examine works from top security conferences on software signing, the design of software signing primitives [53, 52], and empirical studies on software signing (all cited in (section 2)). Our literature review informed us of existing knowledge gaps and provided insights into the design of modern software signing tools. This helped us avoid duplicating questions already addressed in the literature.
  3. 3.
    We designed our questions to align with our formative usability framework (section 3). We developed these questions to focus on identifying the factors responsible for the adoption of Sigstore in practice. Categories D3 and D4 focused on participants’ experiences with their current tools, aligning with the framework’s emphasis on assessing present tooling experiences. Meanwhile, categories D1, D2, and D5 captured participants’ experiences with previous software signing tools, offering insights into their transitions and comparisons over time.
Next, we refined our interview protocols through a series of practice and pilot interviews. We first conducted two practice interviews with the secondary authors of this work, followed by pilot interviews with the first two interview subjects.
Our full interview instrument consisted of three main topics (B-D). As suggested by the table, topics B and C do not fit into the same analysis framework as topic D. Thus, following the guidance of Voils et al. [88], we analyzed those topics in another work [37]. In this work, we focus on the data for Topic D. Topic D consisted of five main question categories, conducted in a semi-structured manner. A summary of the questions in our interview protocol is shown in Table 1.
Table 1: Summary of interview protocol. Topic B and C are analyzed in [37]. We analyze only topic D in this paper. The decision to analyze these parts of the protocol separately was made in accordance with best practices. We present our full protocol in  Appendix A
Topic(# Questions) Sample Questions
A. Demographics (4) What is your role in your team?
B. Perceived software supply chain risks (4) Can you describe any specific software supply chain attacks (e.g., , incidents with 3rd-party dependencies, code contributors, OSS) your team has encountered?
C. Mitigating risks with signing (8) How does the team use software signing to protect its source code (what parts of the process is signing required)?
D. Signing tool adoption (5)
Question Categories in Topic D
D1. What factors did the team consider before adopting this tool/method (Sigstore) over others?
D2. What was the team’s previous signing practice/tool before the introduction of Sigstore?
D3. How does your team implement Sigstore (which components of Sigstore does the team mostly use)?
D4. Have you encountered any challenge(s) using this tool of choice?
D5. Have you/your team considered switching Sigstore for another tool?
4.2 Data Collection
Recruitment: Our study investigates signing tools using a case study of the Sigstore ecosystem. As such, our target participant pool is centered on Sigstore users. Furthermore, our research questions, as reflected in our interview instrument, aim to understand the reasons prior to adoption that teams choose software tools (Sigstore), how Sigstore is implemented and the challenges and benefits of Sigstore.
To effectively answer these questions related to decision-making, we required a target population of high-ranking or experienced practitioners who either oversee Software signing compliance or manage the infrastructure of their respective teams or organizations.
To recruit this target population, we employed a non-probability purposive and snowball sampling approach  [7]. We leveraged the personal network of one of the authors to send an initial invitation to members of the 2023 KubeCon (hosted by the Linux Foundation) organizing committee, who were also part of the In-toto [83, 32] Steering Committee (ITSC). This initial invitation yielded six (6) participants. We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (five(5)) and authors contacts (seven) to recruit an additional twelve(12) participants. We interviewed 18 subjects, but only responses from 13 participants fit our analysis framework [^1] We interviewed 18 participants but ultimately excluded data from five for this study. Two of them used software signing tools other than Sigstore, and three gave limited information about their software signing tool..
Participant Demographics:
Table 2: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff. We highlight type of software subject’s immediate team produces. * refers to cases where subject’s team built other types of software in addition to the main stated. O refers to cases where subject belonged to multiple teams with other products to the one stated. We also highlight which participants were recruited from the In-toto Steering committee (I) and which were from snowball sampling (S).
ID Role Experience Software Type Rec.
P1 Research leader 5 years Internal POC software I
P2 Senior mgmt. 15 years SAAS security tool*O S
P3 Senior mgmt. 13 years SAAS security tool*O I
P4 Technical leader 20 years Open-source tooling S
P5 Engineer 2 years Internal security tooling S
P6 Technical leader 27 years Internal security tooling*O S
P7 Manager 6 years Security tooling*O S
P8 Engineer 2.5 years SAAS security* I
P9 Technical leader 4 years Consultancy S
P10 Research leader 13 years POC security Software* S
P11 Senior mgmt. 15 years Internal security tooling I
P12 Senior mgmt. 15 years SAAS S
P13 Manager 11 years Security tooling S
We summarize our interview participants’ demographic information in Table 2. Our interview sampling approach yielded 13 experienced security practitioners who indicated that they are responsible for initiating or implementing their organization’s security controls and (or) have compliance-related responsibilities. This involvement gives them the relevant context to assess their organization’s strategies for adopting software signing tools.   Interviews: We conducted our interviews over Zoom. Each interview lasted a median of 50 minutes. The lead author conducted these interviews. We offered each subject a \$100 gift card as an incentive, in recognition of their experience level.
4.3 Data Analysis
We transcribed our interview recordings using www.rev.com’s human transcription service and anonymized all potentially identifiable information.
To analyze our interviews, we employed a thematic analysis approach [9] alongside our modified usability framework (section 3, Figure 1) within an interpretivist paradigm [4].
An interpretivist approach allows each participant’s response to be treated as a unique case, considering its specific context and surrounding variables. This is particularly valuable in our study, as participants belong to different teams with diverse implementation contexts and varying uses of Sigstore.
Our analysis began with a data familiarization stage [80] and proceeded through three phases, carried out by two analysts  [^2] One of the analysts is the lead author.. In the first two phases, we inductively memoed and coded our transcripts, developing our initial codebooks without referencing our usability framework [21]. In the final phase, we derived themes from our inductively generated codes and then applied our usability framework to further refine the extracted themes. We describe these steps in detail next.
Memoing & Codebook Creation: First, beginning with the initial 18 transcripts, the two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations and techniques outlined by O’Connor & Joffe [56] and Campbell et al. [12]. This approach ensured the reliability of our analysis process while optimizing data resources. [^3] Although there is no consensus on the ideal proportion of data to use, O’Connor and Joffe suggest that selecting 10–25% is typical [56]. We randomly selected approximately 33% (six transcripts) to develop our codebook. This slightly exceeded O’Connor & Joffe’s recommended percentage to allow the analysts a better familiarization with the nature of our collected data, and analysis format.
After each round, emerging codes were discussed in meetings, resulting in a total of 506 coded memos.
We then reviewed the coded memos and finalized an initial codebook with 36 code categories. To assess the reliability of the coding process, we combined the techniques outlined by Maxam & Davis [51] and Campbell et al. [12]. Specifically, we randomly selected an unanalyzed transcript, which we coded independently. Following Feng et al. ’s [20] recommendation, we used the percentage agreement measure to evaluate consistency, as both analysts had contributed to developing the codebook. We obtained an 89% agreement score and resolved all disagreements through discussions and code refinements. This high agreement indicates the stability of our coding scheme, allowing us to reliably singly code the remaining transcripts [12, 56].
Codebook Revision: Following the reliability assessment of our codebook, and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six transcripts per round. In the first batch, six new code categories were introduced, which were subsequently reviewed by the other analyst. The same analytical process was applied to the final six transcripts, leading to the addition of four new code categories, which were subsequently reviewed by the other analyst. Notably, of the 10 codes added during this phase of codebook refinement, only three provided new insights into the data. By mutual agreement between both analysts, these were categorized as minor codes.
To ensure our sample size and analysis was adequate, we calculated code saturation for each category in our codebook. Following the recommendations and methods of Guest et al. [26] and Maxam & Davis [51], we measured the cumulative number of new codes introduced in each interview. The results are presented in Figure 6. As shown, saturation was achieved after interview 11 , with no new codes emerging beyond this point. Each interview had a median of 15 unique codes (for topic D alone), as illustrated in Figure 6.
Thematic Analysis & Usability Framework Analysis: Next, following the method proposed by Braun & Clarke [9], we derived initial themes from our code categories. We focused on identifying themes related to the D part of our protocol, as highlighted earlier in Table 1. We then conducted another round of thematic analysis, incorporating the initial themes alongside our formative usability framework (section 3). In this phase, we mapped our initial themes onto the factors defined within our framework.
Survey Attempt: In addition to our interviews, we designed and conducted a survey to gather insights from members of the Sigstore community who might not have had the time for an interview. However, our survey received only a few responses, highlighting the challenges of conducting surveys within highly specialized communities [78]. We elaborate on our survey results and details in Appendix D.
4.4 Study Limitation & Threats to Validity
To discuss the limitations of our work, we follow Verdecchia et al.’s [87] recommendation to reflect on the greatest threats to our research. In addition, we discuss how our experiences and perspectives may have influenced our research.
Construct Validity: Cresswell’s usability framework was developed for health information technology. We adapted it to software signing tooling, but may have introduced flaws.
Internal Validity: Threats to internal validity reduce the reliability of conclusions of cause-and-effect relationships. An additional internal threat exists in the potential subjectivity of the development of our codebook. To mitigate this subjectivity, we utilized multiple raters and measured their agreement.
External Validity: The primary bias to our work is that we recruited from a population biased toward the use of Sigstore. We viewed this as necessary in order to recruit sufficient subjects, and feel that that result — a focused treatment of usability in Sigstore — is a valuable contribution to the literature. A secondary threat arises from the sample size of our work, which is due to the challenges of recruiting expert subjects. A final threat to the validity of our research is that it reflects the state of Sigstore at only one point in time. Our interviews were conducted between November 1, 2023, and February 9, 2024 and reflect the experiences of Sigstore users at that time.
Statement of Positionality: We acknowledge that one of our paper’s authors is a Sigstore tool contributor. This author was not involved in developing the research design nor the initial analysis of the results, but provided insight in later analysis.
5 Results
5.1 RQ1: Problems & Strengths of Sigstore
Table 3: Summary of Practitioner-Reported Advantages of Using Sigstore
Topics & Associated Examples Subjects
Technological Factors
Ease of Use 8
Signing Workflow & Verification P1, P2, P7, P8, P10-13
Setting up with automated CI/CD actions P9
No key distribution problems P1
Use of Short-lived Keys & Certificate 3 –P2, P3, P11
Signer ID Management 4
Use of OIDC(Keyless) to authenticate signers P3, P5, P10, P12
Compatibility with Several New Technologies 4
Integrability with SLSA build P12
Integrability with several container registries/technologies P9, P10
Integrability with several cloud-native applications P11
Precence of a Transparency Log 3
Transparency logs increase security P5, P10
Evaluation of signing adoption using logs P8
Bundling Signatures With Provenance Attestations 2 – P3, P4
Reliability of Service 1 – P7
Macroenvironmental Factors
Free/Open-Source 2 – P7, P13
The first aim of our research work was to ascertain from the the experiences of our subjects the problem and strengths they have experienced while using Sigstore. We categorized each identified theme based on our usability framework factors: Technological Factors – situations where observed conditions were linked to the behavior of Sigstore technology itself, Human/Social Factors – participants’ observations related to gaps in support resources, previous experiences (with other tools) Organizational Factors – aspects tied to users’ organizations or their organizational capacities, and Macroenvironmental Factors – external influences affecting adoption and use.
5.1.1 Sigstore’s Strength
We present a summarized list of practitioner-reported advantages of using Sigstore in Table 3. First, we extracted the benefits our participants experienced while using Sigstore from their responses. These reflections capture their firsthand experiences with Sigstore.
We grouped their responses into eight categories: ease of use, identity management for signers, compatibility with emerging technologies, use of short-lived keys and certificates, implementation of a transparency log, encouragement of bundling signatures with attestations, free access to the public Sigstore instance, and reliability of the Sigstore service. When mapped to our modified usability framework, these factors fell under two categories: Technological and Macroenvironmental factors from our framework.
Technological Factors
Ease of use: Ease of Use was a predominant reason why many participants chose to use Sigstore, making it the most frequently cited advantage over other Software signing tools. Participants commonly highlighted the simplicity of following the Sigstore workflow for signing and verifying signatures. P7 (security tools, cloud security) states, “ I don’t really know of any other options that provide the same conveniences that Sigstore provides… It really is just one command to sign something and then one more command to verify it, and so much of the work is handled by Sigstore.” Other advantages associated with Ease of Use, as mentioned by practitioners, include the seamless integration of the Sigstore workflow into CI/CD automation using GitHub Actions and the reduced complexity of key management.
Identity management for Signers: Sigstore’s integration of OIDC identity management was a key strength reported by our participants. In traditional software signing, signer identity is often an ad-hoc feature that may or may not be required by the signing authority. In Sigstore, however, this feature—implemented via OIDC—is integrated with the Fulcio certificate authority as an identity management step before creating a signature. Participants greatly valued the elimination of the key generation and distribution phase, as it reduced the time spent distributing keys to key servers and searching for keys, streamlining the signing process. P12 (SAAS, SSC security tool) Highlights this point, “The other strength that it has is I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key … that’s super useful because I could say, "Oh, this is signed with Mike’s public GitHub identity." So unless my GitHub identity has been compromised, that’s much better.”
Compatibility with Several New Technologies: A portion of our participants reported that Sigstore’s integration and compatibility with a wide range of modern technologies provided a significant advantage. The mentioned technologies included security tooling such as SLSA build attestations, as well as containerization technologies and cloud computing applications. In most cases, the resources required to set up connections to these technologies were minimal or nonexistent. P9 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers pretty trivial to do if you use something like cosign…Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is pretty easy ”
Use of Short-lived Keys/Certificates: Sigstore’s enforcement of short-lived keys and certificates as a security feature was recognized by our participants as a major strength of the tooling. Traditional software signing tools require users to manually set expiration times for cryptographic elements such as signatures, keys, and certificates at the time of generation. While this approach provides customization flexibility, it also introduces the risk of long-lived cryptographic materials remaining valid for extended periods, increasing the likelihood of compromise. Participants appreciated Sigstore’s automated management of key and certificate lifetimes, which mitigates this risk. P2 (SAAS, SSC security tool) states, “So (with) PGP, you still have to figure out the key distribution problem. Because Sigstore uses short-lived keys that solves a lot of those issues around key storage and key distribution…”
Presence of a Transparency Log: The inclusion of a transparency log in the Sigstore ecosystem, serving as a tamper-resistant secondary record of changes to an artifact’s metadata, was seen as highly beneficial by participants. This feature provided an additional layer of trust, allowing users to verify signed artifacts with greater confidence. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P8 (SAAS, SSC security): “From a higher level, we’re going to use a transparency log to determine what was actually signed and what was written into that transparency log.”
Other technological factors noted by participants are the reliability of Sigstore’s public instance was a valuable aspect of Sigstore.
Macroenvironmental factors
Free/Open-Source:
Sigstore’s open-source nature was also highlighted by participants as a key advantage. They noted that the no-cost solution reduces both setup and maintenance barriers, particularly for users leveraging Sigstore’s public deployments. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public”
5.1.2 Sigstore’s Weaknessses
We summarize the difficulties reported by participants in using Sigstore in  Table 4. Most of the weaknesses reported often belong to multi-factors. e.g., In the problems associated with Enterprise applicabiliirty of Sigstore, Some participants reported this as a consequence of regulations place on organizations which are both macroenvironmental (regulations) and Organizational.
Table 4: Reported Difficulties Encountered by Participants in Using Sigstore. We indicate the associated usability factor of each category of weakness using – T-technology, P-social/human, O-organizational, M-macroenvironmental factors.
Topics & Associated Examples Subjects
Enterprise Adoption Limitations 7
Rate Limiting Problems – T P3, P7, P10, P11
Lack of dedicated Support & Maintenance – T & P P11, P2
Not Suited for Regulated Organizations – M & O P3
Latency Concerns – T P6
Transparency Log Issues 6
Not Suitable for private Setup – T P2, P3, P6, P10, P13
Use in Air Gap Conditions – T P2, P3, P8
Efforts to Monitor Logs – P P2
Private Sigstore Instance Setup 5
Documentation – P P8, P9, P6
Limited Community Support – M P11
Infrastructure Requirements & Maintenance Costs – T P5, P6
Other Documentations and Usage Information IssuesP 3 — P1, P10, P13
Integration to Other Systems 3
Attestation Storage – T P1
Gitlab & Jenkins – T P8
Other Unsupported technologies – T P12
Offline CapabilitiesT 2 – P3, P4
Fulcio Issues 1 – P3
Timestamping Issues – T
Fulcio-OIDC Workflow – T
Software LibrariesT 1 – P7
Enterprise Adoption Limitations: Many of our participants mentioned that Sigstore is not suited for large enterprise applications. The issues reported by participants include rate limiting—where Sigstore restricts the number of signatures that can be created per unit time—along with latency concerns, where the service’s turnaround time for a large volume of signatures is problematic. Additionally, participants highlighted the lack of dedicated support and maintenance teams due to Sigstore ’s open-source nature, as well as its unsuitability for organizations in regulated sectors for the same reason. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P8 (SAAS, SSC security): “I think the major challenge was that we were relying on the public Sigstore instance, and I don’t think it could meet our signing needs in terms of just the capacity of signatures we needed as an enterprise. We would constantly be getting rate limited by the service.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “For customers that are very large in scale, the biggest of the Fortune 100 or customers operating in very highly regulated environments, I think operating your own Sigstore instance within that air gap in a private environment could be valuable. And I think it depends on the amount of software you’re producing and the frequency with which you’re producing it.” The issues here were a mixture of all usability factors.
Transparency Log Issues: While the transparency log bundled with Sigstore was highlighted as a major advantage by participants, it was also a significant concern. Their concerns stemmed from the public nature of the log, where sensitive artifact metadata from company assets could be exposed in a publicly accessible record. Another common issue raised was the log’s usability in air-gapped environments—systems that are physically isolated from unsecured networks for security reasons. Additionally, participants noted the manual effort required to continuously monitor and manage the log. Most of these participants were in the process of experimenting with setting up the transparency log at the time of the interviews. Transparency log issues were mostly technological.
{quoting}
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P6 (security tooling, digital technology): “We’ve got some teams piloting Rekor, for instance, for different traceability usages and so on…I’m all for the transparency log, but we just have to be careful who it’s transparent to at what point in time.”
Setting up Private Sigstore Instance : Sigstore’s customizability was a feature frequently utilized and discussed by participants; however, it was also associated with certain challenges. The primary issues identified included limited documentation on setup, scarce community usage information, and the infrastructure requirements for private deployments. Issues here were primarily related to participants’ need for support resources, which where macroenvironmental and social in nature. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P11 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to help people deploy Fulcio, like on prem, or in their own infrastructure.”
Other Documentation Issues: Overall, the documentation for Sigstore, along with practical usage information and support, was found to be insufficient. Participants noted that updates to the documentation lagged behind changes in the Sigstore product and that there was a lack of clarity regarding the use of different Sigstore components. P13 (security tooling, internet service) highlights this, “ I would say the documentation lacks a bit. Just I think if you look at the documentation, it was updated quite a long time ago, like the ReadMe file. And so if we could have these different options saying that, okay, if you are starting over, this is how you can do rekor and Fulcio.”
Integration to Other Systems: Participants also expressed a desire for a wider range of compatible technologies beyond what Sigstore currently supports. Commonly mentioned integrations included other CI/CD platforms such as Jenkins and GitLab, as well as attestation storage databases. These limitations were often attributed to Sigstore being an emerging technology, as highlighted by P12 (SAAS, SSC security tool), “The weakness is not everything supports it. There’s a lot of weaknesses just around not just it’s being a newer technology”
Fulcio Issues: Some Sigstore issues were also tied to its certificate authority’s (Fulcio) time stamping capabilities, and implementation of OIDC for keyless signing. Fulcio acts as an intermediary using OIDC identities to bind to a short-lived certificate, this was criticized by P3. P3 also commented on the initial lack of time stamping abilities for the initial versions of Fulcio. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “I think we do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider. And then I think the other big thing that we did, because Sigstore originally didn’t support it, was time-stamping. Because we’re using Fulcio, we want to ensure that the signature was created when the certificate was valid. And so we use external time-stamping authority in order to be able to do that.”
Offline Capabilities: Another reported limitation of Sigstore is the absence of offline verification capabilities using offline keys. This issue is related to the transparency log’s inability to function in air-gapped environments. However, in this case, participants specifically desired the ability to use Sigstore’s signing and verification features in offline scenarios.
Software Libraries: The programming libraries of Sigstore were also mentioned as a challenge, particularly in cases where users wanted to integrate Sigstore’s signing capabilities directly into an application.
5.1.3 Effect of Identified Problems on Sigstore components used by Subjects
Table 5: Sigstore Functionalities Used by Subjects
Tool # of Subjects
Cosign (Signing & Verification CLI) 9
Certificate Authority (Fulcio) 9
Keyless Signing & Identity Manager (OIDC) 9
Transparency Log (Rekor) 8
Gitsign (Commit Signing) 4
Component Customization & Local Sigstore Deployment 2
We assessed the effect of participants’ reported problems with Sigstore on their usage of each component. Table 5 summarizes our subjects’ reported usage of Sigstore components and functionalities.
Key components like Cosign, Fulcio, and keyless signing saw high adoption, while local Sigstore deployment had low usage due to numerous reported issues. Several participants cited difficulties with documentation and setup, and as a result, only two successfully deployed private instances, likely due to these challenges (Table 4, section 5.1.2).
The Rekor transparency log presents a mixed case, with both high adoption and significant reported issues. Of the eight participants who used it, three were still in the pilot phase. Although Rekor was recognized as a key factor driving adoption due to its strengths, it also faced a high number of reported issues, indicating that, despite its benefits, challenges persist for some participants.
Other Components Without Mention Though the participants in the survey shared experiences with user-facing components, discussion about other crucial elements was notably absent. In particular, there was very little mention of log witnessing [28], or monitoring [22] for either Fulcio or Rekor. This is remarkable, as transparency solutions require the existence of these parties to ensure log integrity and identify log misbehavior. We believe that targeted future work should identify possible gaps in user perception, as well as identify user stories so users verify log witnesses and monitor information.
5.2 RQ2: How and Why Practitioners Change Software Signing Tools
Table 6: Reasons Practitioners Choose or Switch to Sigstore Before Adoption.
Topics & Associated Examples Subjects
Human/Social Factors
Practitioners Contribute to Sigstore 6 – P2, P4, P6, P7, P8, P10
GPG Issues 6
Low adoption rates P1, P10
Key management issues & Other usability concerns P1, P6, P9, P12, P13
Steep learning curve P9, P12
Compatibility with newer technology P12
Notary Issues 2
Non-demand from customers P9
Compatibility with other tools P9
Lack of regular updates P9
Key & Identity Management P5
Proprietary Tool Issues 1
Difficult to setup P11
Technological Factors
Available Sigstore Functionalities 3 – P1, P5, P10
A transparency log, etc P5, P10
Compatibility to other Tools P1
Macroenvironmental Factors
Regulation & Standards 4 – P5, P6, P11, P13
Large User Community 1 – P8
Inherent Trust of Creators 1
Trust of CNCF products P3
Figure 3: Practitioners’ Evolution of Software Signing Tools. Most reported their organizations moved from no former tools, a form of GPG signing or notary to Sigstore. Most participants depicted with the flow from none to a sigstore implementation report that; they either initially used no tools, or have always used Sigstore.
We also tackle the usability of software signing tool especially as it relates to decisions that prompts the switch of practitioners in using different tools. In the context of our study, we asked participants what factors prior to adopting Sigstore influence practitioner’s to switch to Sigstore. The factors here are mostly Macroenvironmental (relating to regulations, user communities), Human (Previous experiences), and Technological (available capabilities of sigstore)
First, we present a Sankey flow diagram ( Figure 3) illustrating the evolution of signing tool usage among our participants’ teams up to the present. A significant portion of participants (6), as depicted transitioned from using no tools to adopting Sigstore —either they had always used Sigstore since the inception of their organization or continued using their initial signing tools alongside Sigstore. The most notable transition was from GPG-based signing implementations (3) to Sigstore, followed by a shift from Notary (2). Other reported transitions included Skopeo (1), unspecified internal tools (1), and a proprietary tool (1).
The reasons most of our participants chose to switch to Sigstore stemmed from a combination of two key factors: issues related to a subpar user experience with previous tools and advantages associated with Sigstore. We summarize these factors in Table 6.
Contributions to Sigstore: Being part of the Sigstore community as contributors was a major reason for adopting Sigstore. Some participants (e.g., P6) stated that their organizations have a policy of contributing to open-source communities they intend to adopt. Others noted that their organizations were relatively new compared to Sigstore—meaning that Sigstore had already existed before these organizations were established, and some of their members had contributed to Sigstore before joining their respective organizations. Additionally, some participants mentioned that key members of their organizations were directly involved in the initial development of the Sigstore project. P10 (POC software & security tool, cloud security) reflects this, “I’d say many of the early staff, the first 10 to 20 staff were involved in the Sigstore community. There’s not just our interest as a company in a Sigstore, but there’s a lot of personal ties to Sigstore.”
Available Sigstore Functionalities: Sigstore’s advertised functionalities also influenced its adoption. While several participants were drawn to the transparency log (as highlighted in section 5.1.1), its integrability with existing technologies they were using at the time also played a role in their decision. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P5 (internal security tool, cloud security): “… Sigstore has other features, like you can use your OIDC identities, like your GitHub identities as well, to do the signing, as well as they also maintain a transparency log as well …” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P1 (POC software, digital technology): “Factors I considered were, …, the existence of tools to use it.”
Regulations and Standards: The impact of regulations and standards on adopting Sigstore was significant among our participants. Some participants who previously used other tools (2) switched to Sigstore due to security frameworks and standards recommending it as a best practice. Additionally, the remaining participants (2) noted that regulations played a partial role in their decision to adopt Sigstore.
{quoting}
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P5 (internal security tool, cloud security): “…standards actually really helped to convince everyone that we should be doing this … But once a standard is put in place, once the requirements are set, then it just sped up the process of adopting these new technologies like Sigstore …”
User Community & Trust Of Sigstore Creators: The large user community of Sigstore was also a key reason for its adoption among participants. This aligns with feedback from some participants (P1, P10) who cited GPG’s smaller user base as a drawback. Additionally, some participants expressed inherent trust in the CNCF (Cloud Native Computing Foundation) security tooling community, which they had grown to rely on over time. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “We do defer often at the higher level projects to different foundations. So we would be more likely to trust something from the CNCF, the Cloud Native Computing Foundation or Sigstore rather than small little independent developer projects or things like that.”
GPG & Notary Issues: Issues associated with previously existing tools played a major role in practitioners’ decisions to switch to Sigstore. GPG-based signing implementations have historically been criticized for key management challenges and usability concerns. Our participants echoed these issues, citing additional factors such as low adoption rates, steep learning curves, and limited integration with modern technologies.
Notary-related concerns primarily revolved around customer demands, compatibility with other technologies, and infrequent updates. Key and identity management challenges were common to both GPG and Notary. Meanwhile, users of proprietary signing tools faced difficulties with complex setup processes.
6 Discussion
In this section, we highlight two reflections from our data.
The role of community in software signing tool adoption: A substantial number of participants reported that they adopted Sigstore because of their involvement as contributors to the Sigstore tooling community (section 5.2). This reflects a key macroenvironmental factor from our usability frameworks that influenced users’ decisions to adopt these tools. Similarly, several reported challenges were linked to the lack of community support in setting up private Sigstore instances and accessing usage information (section 5.1.2).
These findings highlight the critical role of community, as emphasized in prior research [72]. New adopters often depend on communities they trust, contribute to, or receive support from, which significantly shapes their decision to adopt such tools. Toolmakers should prioritize fostering and maintaining a strong community around their tools, as this can greatly influence adoption and success.
Privacy Concerns: Privacy concerns emerged as a big issue among users of Sigstore’s Rekor transparency log module, with 5 out of 6 participants who reported drawbacks citing privacy as a significant concern. This issue also affects the use of OIDC for keyless signing. While this challenge is not unique to Sigstore [52] and is a recurring issue in software signing and modern cryptography, it underscores the need for continued innovation to address privacy-preserving mechanisms that balance transparency with user privacy, ensuring broader adoption and trust in cryptographic tools like Sigstore. While works like Speranza[52] has been proposed, efforts should focus on integrating these solutions to tools like Sigstore.
7 Conclusion
Software signing allows software provenance to be guaranteed through cryptographic primitives. This property would address many software supply chain attacks. However, for decades, the widespread adoption of software signing has eluded the software industry. In this study, we examined a recent change in software signing practices by focusing on Sigstore, the leading next-generation software signing platform. We interviewed industry experts to understand Sigstore from a usability perspective. We identified the factors they weigh and the challenges they face. Our results shed light on usability issues for software signing in particular, and will inform more usable software tools writ broad.
8 Acknowledgments
We thank the study participants for contributing their time to this work. We acknowledge support from Google, Cisco, and NSF #2229703. We also thank Hayden Blauzvern and all reviewers who provided feedback on this work.
Appendix A Interview Protocol
Table 1 gave a summary of the interview protocol. Here we describe the full protocol. We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style. Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
We include Section D of the protocol for completeness, although it was not analyzed in this study.
Table 7: Our full interview protocol. The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions. We only analyze parts A and D in this study
Themes Sub-themes Questions
A: Demographic A-1: What best describes your role in your team? (Security engineer, Infrastructure, software engineer, etc)
A-2: What is your seniority level? How many years of experience?
A-3: What is the team size?
A-4: $\dagger$ What are the team’s major software products/artifacts?
A-5: What type of Organization/company?
B: Software Supply Chain
Failure Experienced
B-1: Describe briefly the team’s process from project conception to product release and maintenance – This is to understand the unique context of each practitioner’s software production case.
B-2: $\star$ What do you consider a Software Supply chain attack/incident to be?
B-3: Can you describe any specific software supply chain risks (or incidences with third-party dependencies, code contributors, open source, etc.) that your team has encountered during your software development process? How were these addressed? (Alternatively – How did this affect the decision to implement software signing?
[…]
B-5: What is your team’s greatest source of software supply chain security risk? Project components(Third-party dependencies, build process, code contributors, etc) $\dagger$ Between the Software Engineering Process vs project components which constitute a greater source of risks?
C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from Section B): why did the team choose to implement software signing? Is software signing the team’s major strategy to secure its supply chain? Any other complimentary security efforts and methods? Are these strategies (software signing and complementary security techniques) influenced by regulations and standards? What challenges or obstacles did your team face while integrating Software Signing into your supply chain security practices? $\star$ How do different team members contribute to the implementation of Software Signing? What roles and responsibilities are involved? (After Practise). Possible Follow up – Do you think Software Signing on its merit (if implemented right) is good enough to completely secure a supply chain? (Added after Pilot).
Third-party dependencies C-2: What are the team’s peculiar selection strategies for Third-party dependencies? How does the presence of a signature influence this decision? How is the authenticity of this signature verified? Any other methods/practices to ensure the trustworthiness of the third-party dependencies before integrating them into your projects? Possible Follow-up – Can you describe these methods/practices?
C-3: What influences the team to discontinue the use of a package? How does the Signature on the dependency influence this decision?
C-4: How does the team manage its third-party dependencies’ security? How do you maintain awareness of potential vulnerabilities or threats related to third-party components?
First-party source code C-5: Is software signing a requirement for team developers (insider threats) and open-source contributors (if any)? how does the team use software signing to protect its source code (what parts of the process is signing required e.g. Commit signing)?
Build Process C-6: (How) Does the team utilize signing in its build process?
Package Artifact/Deployment C-7: How is Signing used to protect the following from compromise? how Artifact’s build binaries/deployment pipeline Artifact’s repository
C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism? (This question is derived from  [36]).
D: Signing Tool Selection D-1: What software signing tooling does the team use (If not mentioned yet)
D-2: What factors did the team consider before adopting this tooling over others?
D-3: What was the team’s previous signing practice before the introduction of current tooling?
D-4: How does your team implement this tooling (which components of tool does the team use)?
D-5: Have you encountered any challenge(s) using this tool of choice? How have you coped with this limitation(s)? Are there any areas where you believe further enhancements could be made in your current implementation of these methods?
D-6: Have you/your team considered switching this tool for another? What other tools have been considered or are currently being considered?
Appendix B Codebooks
Table 8 describes our codebook, with code definition, and example quote tagged with that code.
Table 8: Excerpts from the final codebook we used for our analysis.
Code Definition Sample Quote
Sigstore Criticisms Criticisms and weaknesses associated with implementing Sigstore for signing. S3: "We do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider. "
Sigstore Improvements Suggested improvements for Sigstore. S9: "I would say probably integrations within different platforms, so say you use GitLab or GitHub or Jenkins or something like that. Try to leverage some type of integration to make that, I guess, less friction for the developers to implement, something like that."
Sigstore Strengths Strengths of Sigstore and reasons why subjects prefer using Sigstore. S7: "The convenience and the good user experience is probably what keeps us using Sigstore "
Sigstore Alternatives Considered Alternatives of Sigstore that were or are being considered by the subject (for subjects whose primary signing implementation is Sigstore). S5: "We are also looking at other new alternatives that are coming up, like OpenPubkey. Docker just released a new standard called OpenPubKey, which is trying to solve this issue where they will not require you to have infrastructures and stuff."
Sigstore Components Used Components of Sigstore (Cosign, Fulcio, etc.) being used by the subject. S2: "Mostly fulcio is what’s integrated into our product."
Notary Criticism Criticisms and weaknesses Associated with Notary – promoting users to stop using them. "S5: "We were using Notary, and Notary only verifies that the image was signed using a specific key, but you cannot verify that who is the owner of the key."
Appendix C Traditional vs Next-Gen Software Signing Workflow
We show a typical traditional package signing workflow in Figure 4 compared to a next-generation workflow in Figure 5.
Figure 4: Typical workflow for software signing and verifying signatures. The component author packages (A) and signs (B) their software. The signed package (C) and public key (D) are published. To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
Figure 5: Typical workflow for next-generation software signing. The component author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider. The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.
Traditional Signing: The signature creation process begins when the maintainer’s artifact is ready for submission. The signer generates a key pair that provides them with a private key for signing their software artifact and a public key for a verifier to validate the artifact’s signature. Once the artifact is signed, it must be submitted to a package registry (such as PyPi, npm, or Maven) or the intended party. During this phase, the signer is responsible for making their public key available and discoverable so future users can verify the signer’s identity. In the verification phase, the verifier retrieves the signer’s public key, uses it to decrypt the signature file, and checks for equivalency with the software artifact’s hashed digest.
Next-generation Signing:: Keyless/Ephemeral software signing mitigates the security risk of private key compromise. Keyless software signing differs from traditional public-key methodology in the fact that the signer does not rely on a long-term private key to sign their messages. Instead of the signer being responsible for the secure maintenance of a private key, in keyless signing, the signer uses a short-lived key that only lasts long enough to sign their artifact [2].
Appendix D Survey Attempt
In addition to our interviews, we developed a survey distributed to Sigstore users at the 2024 SigstoreCon conference and through post-conference communications with the Sigstore user community. This aimed to gather perspectives from those who might not have had time for an interview. We exclude this data from our work due to several factors that compromise the quality of the results; these factors include respondents who completed the survey in an unrealistically short amount of time and those who failed to complete the entire survey.
Here, we summarize the answers to an excerpt of the questions responded to in our survey:
What signing tool did your team use prior to the introduction of Sigstore?: Among 7 participants who utilize Sigstore, 3/7 responded that they use PGP/GPG-based signing and 4/7 did not use any toolings prior to using Sigstore.
Which of the following factors have influenced your team’s use of signing? (select all): Among 7 participants who utilize Sigstore, 4/7 selected "Organizational policies," 4/7 with "Best Practices/Other standards and recommendations" and 3/7 selected "Awareness of prominent SSC attack."
Which Sigstore components does your team use? (select all): Among 7 participants who utilize Sigstore, 3 responded to this question, and all 3 participants selected Cosign, Fulcio, and Rekor.
In which of the following ways have you implemented Sigstore?: Among 7 participants who utilize Sigstore, 3 responded to this question, with 2 responding that they use a private deployment of Sigstore and 1 responding that they use a public instance of Sigstore.
Do you use or require Sigstore (signing) at any of the following stages of your product?: Among 7 participants who utilize Sigstore, 3 responded to this question. All 3 participants shared the "Build/Deployment/final product" stage as part of their answer.
How do you verify Sigstore signatures? (select all): Among 7 participants who utilize Sigstore, 3 responded to this question. All 3 participants shared selecting certificate Verification as part of their answer. 2/3 selected that they check Rekor logs.
Appendix E Code Saturation
Figure 6: Saturation curve. Interviews are plotted in the order in which they were conducted. Each interview covered many topics in detail (orange line, blue line). However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
We present a saturation analysis of our studies in  Figure 6. The green trendline (saturation curve) was measured by identifying the cumulative unique codes present in each interview. We observe saturation at the eleventh interview, indicating that no new unique codes were identified after participant 11. Additionally, the blue trendline (number of codes) shows that the lowest number of codes was obtained from subject S4. This is also reflected in the orange trendline (unique number of codes).
Appendix F Code Evolution Diagram
Figure 7 describes the evolution of our codebook over six rounds of revisions.
Figure 7: Codebook Evolution: The codebook underwent six rounds of revisions, with changes highlighted by different colors: round 2 (pink), round 3 (blue), round 4 (green), round 5 (yellow), and round 6 (purple). The final codes from round 6 were used in this paper, and categorized thematically. We combined framework and thematic analyses to develop the codebook. First, we induced codes from each transcript and analyzed them thematically. During the framework analysis, we mapped codes to the four stages/blocks of the software supply chain factory model. Round 1: The codebook was defined based on emerging themes. Round 2: No changes. Round 3: Blue-highlighted codes (5) were added after memoing and coding the second set of transcripts. Round 4: Green-highlighted codes (2) were added. Round 5: No changes. Round 6: Analysts A1 and A2 determined some codes should be further divided. Purple-highlighted codes (4) were added and the codebook was finalized. We focused on the topics of relating to sigstore(signing tools) including, Notary, Vault, Spiffe, GPG, and others.